radomes.openskiesproject.org
GloriaGunzenhause650
GloriaGunzenhause650 GloriaGunzenhause650
  • 02-10-2015
  • Biology
Answered

If you flipped a penny 99 times and each time it came up heads, then the chance of the penny coming up heads again is ____ . 50% 1% 100% 25%

Verified Answer :

Kalahira
Kalahira Kalahira
  • 06-09-2017

The chance of coming up heads 99 times in a row is abysmally small: 50%^99 = 1.57 * 10^-30.
This however has already happened, it does not influence the next penny flip.
A penny has two sides, so the chance of it landing on one side is 1/2 = 0.5 = 50%.
So the result of the next flip has a chance of 50% of coming up heads.

Answer Link

Related Questions

Aan is a type of mass movement that occurs when a thin block of soil rock and debris separates from the underlying bedrock A creep B Landslide CFlow D Eruption
The lot that Dana is buying for her new one story house is 35 yards by 50 yards Danas house plans show that her house will cover 1600 square feet of land What p
What was the profession of einsteins first wife mileva maric
Who went to his grave believing that he had discovered a westward passage to Asia when he actually found the Americas
What sets the coding cards for Kaibot apart from other resources
PLEASE HJELP ME WORTH 10 POINTS IF CORRECT
The caspian sea is connected to Moscow through which waterway
Stephan buys a melon and divides it into 6 servings he eats 13 of the melon over the weekend how many silces of melon does stephan eat over the weekend
3x2x1x1 What is X
Which saharan cities were not a part of any empire